OpenStudy (anonymous):

how many square feet are there in 576 square inches? scale: 1foot= 12inches

OpenStudy (anonymous):

48 inches

OpenStudy (anonymous):

576/12 = 48

OpenStudy (anonymous):

48squared

OpenStudy (mertsj):

Yes. 1 foot = 12 inches and 1 square foot = 12^2=144 square inches. So divide 576 by 144

OpenStudy (whpalmer4):

\[576 {in}^2 * (\frac{1 ft}{12 in})^2 = 576 {in}^2*\frac{1 {ft}^2}{144 {in}^2} =4 {ft}^2 \] Do it like this with the units and you'll make fewer conversion mistakes.